# AI-Driven Predictive Analytics Approach for Early Prognosis of Chronic Kidney Disease Using Ensemble Learning and Explainable AI thanks:

### v1

@@ V-B Example Based Techniques and Implementations
For the example based techniques, Counterfactuals and Contrastive Explanation Models (CEM)s are to be applied to this study. From the Counterfactuals, a specific set of rules would be found out based on which any prediction is occuring. For the Contrastive Explanation Models, local prediction in a test case would be looked at and then the CEM tool would tell us what changes would need to happen for a subset of the selected features for the best performing Random Forest model to predict the contrasting class.
Figure 12: Independent Features and Values for Prediction of Non-CKD (Class 1)
From fig 13, first the prediction of class 1 or the non-CKD prediction is seen based on the independent features. In the same figure the contrasting class prediction is observed where in each row, the Counterfactuals show us the changes required for each feature to predict the contrasting class or class 0 (CKD). In the first row of figure 17, it can seen that “Haemoglobin” feature value has to change from 13.8 of figure 16 to 9.65 in figure 17 for the prediction to be of CKD. In the next rows, it is seen that “Rbc” has to change from 2 to 0 i.e. “Normal” to “Abnormal”. Similarly, “Serum Creatinine (Sc) had to change from 3.07 to 10.9 for the contrasting class (CKD) prediction. So, the counterfactuals tell us what changes are required for each of the features so that the model would predict the contrasting class.
The Contrastive Explanation Models (CEM) explain in terms of model prediction of a particular class and show us the exact contributing features for which it has predicted that class instead of it’s contrasting class. For example, when class 0 or CKD prediction is analyzed, the CEM model gave explanation that Random Forest model predicted 0 instead of 1 because “Pe” is greater than 0.948 and “Hemo” is less than or equal to 12.775 and “Pot” less than or equal to 8.35. Similarly, for the contrasting class in another test case prediction, the model predicted 1 (non-CKD) instead of 0 (CKD) because “Al” is less than or equal to 0.604 and “Bgr” is less than or equal to 127.22.
V-C Interpretability Measures of Ensemble Models in Training
Since primary objective of this research was to apply the Explainable AI tools for prediction and analysis of CKD variables, it is also important to analyze how the Ensemble Models applied here perform on certain measures like Interpretability, Fidelity, Fidelity Interpretability Index (FII) and Fidelity Accuracy Index (FAcc). In order for us to do that, first it was needed to find out the exact number of features that the tree based models deem significant so that the Interpretability of the models can be calculated as per studies conducted at [19]. After calculating the Interpretability, the next calculations of FII and FAcc would get easier.
For our calculation, it is important to set a threshold contribution score of 0.05 for the features. So features having score higher than that are considered important features and rest are masked features or non-contributing featuers. From fig 14 Adaboost classifier, it can be seen that 13 features have higher scores than 0.05.
@@ V-C Interpretability Measures of Ensemble Models in Training
So, (24-13) or 11 features are redundant. So, the Interpretability score as per [20] would be 11/24 or 0.45.
Next, to calculate Fidelity score in the training set, the Accuracy of a model that is equivalent to the original model with explainability and the Training Accuracy of the original Model were needed.
$F=\frac{\mathrm{Acc.ofEquivalentInterpretableModel}}{Acc.ofActualModel}$ (2)
Figure 13: Partial Dependence Plot with Rbcc and Hemo
Figure 14: Feature Importances in Train Set for Adaboost
Figure 15: Feature Importances in Train Set for Random Forest
Figure 16: Feature Importances in Train Set for Xgboost
Here, a Decision Tree as the equivalent interpretable model is being considered for the other ensemble models. The training accuracy was 1.0 for all the ensemble models and for the Decision Tree the training accuracy was 98.03. Hence, the Fidelity scores of all the ensemble models would be 98.03.
The other two metrics in the interpretability are the Fidelity Interpretability Index (FII) and Fidelity Accuracy Index (FAcc).
@@
TABLE XI: Interpretability Measures In the Training Set of Ensemble Models
@@
AdaBoost 0.45 0.98 0.44 0.98
Random Forest 0.70 0.98 0.68 0.98
XgBoost 0.75 0.98 0.73 0.98
So, although the FII measures are varying, the FAcc measures are same for 3 models as the Fidelity (F) is being multiplied with the Accuracy (Acc) and the training accuracy is 1.0 in all cases.
V-D Interpretability Measures of Ensemble Models in Testing
In this case, the ensemble models would be evaluated on the interpretable measures but on the testing set accuracy metrics. For this purpose, the XAI tools would be utilized to find out the redundant features as before. Since the Adaboost model had less score in the interpretability, primarily focus was shifted on finding the four interpretability measures from Random Forest and XgBoost. This is where the SHAP summary of global test unseen data would be utilized.
Figure 17: Feature Importances in Test Set for Random Forest
Figure 18: Feature Importances in Test Set for Xgboost
From figure 21, it can be seen that when a threshold of 0.05 is set as the contribution score, only 5 features are above that range and hence (24-5) or 19 features have been deemed less contributing by the Random Forest model according to the SHAP summary plot. Similarly for XgBoost, only 3 features are above the strong contribution score range and hence 21 features are less contributing. So, table XI would show the interpretability measures in the test set of these ensemble models.
To Calculate the value of Fidelity (F) in the test set, table VI was referred which has the accuracy metrics in the test set for the equivalent Decision Tree and the accuracy measures of the models in question. It can be seen that the Random Forest model with the best set of parameters found from the Grid Search on 5 fold cross validation has less scores in terms of Interpretability and Fidelity Interpretability Index (FII) but it has higher score in the Fidelity Accuracy Index (FAcc). The comparison between these two models show that Random Forest model has identified more features as important features for contribution rather than XgBoost and since Interpretability (I) is a ratio of important and masked features as per [19]. So XgBoost model has higher score in this metric and naturally in the FII index.
TABLE XII: Interpretability Measures In the Test Set of Ensemble Models
Model Name Interpretability Fidelity FII FAcc
Random Forest 0.79 0.98 0.77 0.96
XgBoost 0.87 0.96 0.83 0.95
V-E Comparison with Existing Research
@@
Exisiting research on the applications of Explainable AI techniques have mostly focused on Feature Based Techniques like Shapley Additive Explanations (SHAP). To this end, study conducted by Gabriel et al is a quite comprehensive study where they employed Case Based Reasoning (CBR) approach to identify causation of CKD from case by case basis on Colombian population [21]. Similar to the work conducted [22] on investigation of XAI models on CKD from patient dataset gathered from a local hospital in Taiwan, Local Interpretable Model Agnostic Explanations (LIME) were employed to explain the predictions of the models for each cases of CKD or non-CKD. Most of the XAI feature based tools were used like Explain to me Like I’m a five year old (Eli5), Partial Dependence Plot (PDP) and Accumulated Local Effects (ALE) to validate which features contribute in which particular value range and also to analyze the inter feature contribution scores. Moreover, the inter feature dependence plots were analyzed for global test cases using SHAP.
@@
The more novel feature of this study is the use of Example Based Techniques like CounterFactuals and Contrastive Explanation Models (CEM) to explain what changes needed to happen to each feature for a change in the outcome variable. To the best of our knowledge, no published work so far has performed similar analysis on the tabular dataset of this CKD UCI ML one.
The Fidelity Interpretability measures were also employed to compare our study with that of [20] and found out that the XGradient Boosting (XgBoost) model of ours has similar performance in the Fidelity Interpretability Index (FII) measure to that of their best performing model at [20] and [22]. The FII value might be little lower than theirs as a threshold of 0.05 was specified and this study considered the features which have more contribution than this measure are the most significant ones.
@@
The difference of this research with that of [20] is that this research has eradicated the feature “Specific Gravity (Sg)” from our analysis beforehand after consulting with a nephrologist in Bangladesh where her feedback was, they don’t consider at any stage the Specific Gravity to be an important factor for CKD prediction. Moreover, features like Pus Cell Count (Pcc) has also been removed as this feature is mostly used for determining Acute Kidney Injury (AKI) which is not the current scope of our research. On the data pre-processing stage, the difference of our study with the concurrent ones, lie in analyzing the pattern of missingness of data, imputation through calculation of RMSE values first, validating the statistical significance of features with feature co-relations and employing several techniques for feature elimination. So, to sum up, validating all the decisions through multiple measures in the pre-processing to model application to model performance analysis, direct consultation with multiple nephrologists and modifying the study with their feedback etc. are the new scope of work performed in this study different from the exisiting ones.
@@ VI Discussion on the Results
On the results end, most of the models had very high scores for the precision and recall specially when selected subset of 14 features were provided as input. Which indicates that there is few number of false positives and false negatives detected by all the models. For our best performing model Random Forest (In terms of Accuracy Metrics) that was chosen with best set of parameters from Grid Search and 5 fold cross validation and XGBoost, the recall score is 1.0 which indicates there is no False Negatives detected that is no instance of wrong classification of Non-CKD as CKD was seen. Only 1 instance from the confusion matrix of Random Forest in figure 1 was seen which has been falsely classified as Non-CKD where it actually was a CKD case.
The primary focus of this study is to analyze the Machine Learning models decision making process in terms of feature contribution, range of feature values, peak and constant contributions, reasons of choosing a certain prediction, changes needed from feature ends for a contrasting class prediction etc. To this end, the Interpretability Fidelity measures were employed. Analyzing in terms of accuracy metrics and interpretability measures, the XgBoost model performs better than all other models specially when Fidelity Interpretability Index (FII) is considered. But the Random Forest model’s performance can’t also be ruled out as it has higher accuracy, F1 score and Fidelity Accuracy (FAcc) value and it considers more input features as important features than XgBoost model. So, from a clinical perspective both the models are presentable to the clinicians where the Random Forest model would provide a few more features as important contributors for prediction than XgBoost model.
This study focused more on the interpretability of the models with explainable AI to make the decision-making process of the model easier to understand for the general people. All the techniques like Eli5, LIME, SHAP, ALE have validated the feature contributions, inter-feature dependence, feature values in CKD vs Non-CKD predictions. For this scope of work, both the feature based and example based explainable AI techniques were focused. To our knowledge, no research has yet implemented the example based techniques like Counterfactuals or Contrastive Explanation Models (CEM) for the domain of using Explainable AI in Chronic Kidney Diseases. But study conducted on heart disease analysis at [19] use example based models like Anchors, Counterfactuals, CEM. If the ML models do well in terms of interpretability, it makes the job of the researchers, clinicians and the general people easier on all ends to understand the prediction models.
@@ VII Limitations of Current Dataset and Potential Improvements
From the dataset perspective, this study is dealing with a CKD dataset containing 250 patients with CKD and 150 patients not having CKD. Since the patient distribution is organized like having CKD patients occurring consecutively and similarly non-CKD patients, that’s why the Machine Learning models have performed really well in the binary classification task. It would have been interesting if the CKD vs Non-CKD was more unevenly distributed and the features accounted for less variance in the data. Also, from the Pearson Correlation, it is found very strong correlation coefficients among features which was also validated through statistical measures on the Logit model. Moreover, there was strong positive variation in the correlation accounting for high accuracy scores. These factors have played important roles in the prediction results of the ML models.
@@
When we had our one to one session with two different nephrologists from Anwar Khan Modern Hospital and Asgar Ali Medical College (Dr. Md. Nabiul Hassan Rana) in Dhaka, Bangladesh, both of them informed us that the current dataset needs more variables for a conclusive prediction of the CKD patients. Based on the class denotion in the dataset, some of the records does not really contain conclusive reasons for declaring a patient having CKD. For example, a 24 year old patient with only history of Diabetes Mellitus had been declared having CKD without any other discriminative feature values. It is unusal for a patient of 24 years to develop CKD so that’s where their suspicion arose. Moreover, another patient with high Albumin presence was declared having CKD without any other conclusive feature values. The high protein in the urea of that patient could happen for a multitude of reasons like heavy exercise, steroid, excess protein intake etc. Some of the features like ‘Pcc’ is used much later to track the progression of CKD to AKI, not to predict the cases of CKD. Features like “Sg”, “Bgr” are not important to predict the cases of CKD. The large percentage of missing data in a lot of features is also a critical aspect to this dataset. Caution was applied in imputation of these missing values for a critical dataset on CKD with patient vitals after analyzing the pattern of missingness and then using RMSE error metrics to compare the performances before and after imputation. To address the dataset issues we tried to get in touch with the contributor of the dataset in Apollo Hospital, Medurai India, but to no avail.
VII-A Approach to developing a new dataset

### v2

@@ V-B Example Based Techniques and Implementations
For the example based techniques, Counterfactuals and Contrastive Explanation Models (CEM)s are to be applied to this study. From the Counterfactuals, a specific set of rules would be found out based on the prediction. For the Contrastive Explanation Models, local prediction in a test case would be looked at and then the CEM tool would tell us what changes would need to happen for a subset of the selected features for the best performing Random Forest model to predict the contrasting class.
Figure 12: Counter Factual for the Original Prediction of ‘Non-CKD’ versus the contrasting ‘CKD’ predictions
From fig 12, the last row “Original” represents the independent feature values for a Non-CKD prediction for a 24 year old patient. The other 5 counterfactuals in the rows above indicate the corresponding changes for each feature value for the contrasting class ‘CKD’ prediction. It is seen from counterfactual 1 and 3 that high value in ‘Sc’ or Serum Creatinine is present for CKD prediction. Also, the ‘Albumin (Al)’ quantity is high for the 2nd case of CKD prediction. But if we focus on the last row ‘Original’ prediction of Non-CKD we can observe that, in order for the model to predict the Non-CKD case, the independent feature value for ‘Al’ is very low and ‘Sc’ which is the fundamental bio-indicator as clinical marker for CKD are also very low. So if this heatmap is presented to a physician, the case of CKD examples could be presented and informed that, to prevent the CKD from progressing to the next stages these are the factors that need to be further investigated upon.
V-C Interpretability Measures of Ensemble Models
Since primary objective of this research was to apply the Explainable AI tools for prediction and analysis of CKD variables, it is also important to analyze how the Ensemble Models applied here perform on certain measures like Interpretability, Fidelity, Fidelity Interpretability Index (FII) and Fidelity Accuracy Index (FAcc). FII signifies the combined score of the model in both fidelity and interpretability. The higher the score, the more trust we can put in the model for it’s decision results and predictions. Also, we can explain the decisions in an easier way if the FII score is higher for a model. FAcc signifies a balanced score of Fidelity with the Actual Accuracy of the model. Since we want both accurate model that we could trust, that is why first of all we need the model predictions to be accurate and then look for trust in the models decision process. First we would calculate the Interpretability (I). Studies conducted by Thomas Norrenbrock et al at [67] have established that the fewer the number of features chosen while evaluation, the easier it is to interpret the model decision-making to the laymen. To find the interpretability score, we would first find out the feature importances in the test set of the best fold in the 10 fold cross validation of the Random Forest, Adaboost and Xgboost model. We would also employ SHAP as a surrogate model to find out feature importance scores across the global test cases. Then use a similarity measure i.e. Cosine Similarity to find out how similar the feature contribution scores are for the two different models. The cosine similarity in equation (3) is specified for our use case by changing the original ‘u’ and ‘v’ variables. The reason for us to use this similarity metric is, we don’t just want to rely on ensemble models feature importance, rather we want to validate with SHAP for correctly estimating number of important vs. redundant features. To find a threshold for feature importance, we would rather employ the cumulative feature importance scores and discard features that contribute to less than 10% on the cumulative feature importance scores. For the cosine similarity in equation (3), EM represents the feature importance score from the model and LIME represents the feature importance score by LIME explanations.
@@ V-C Interpretability Measures of Ensemble Models
Looking at the feature importance scores of the Xgboost in figure 13 and figure 15 by LIME, it is observed that 4 features are contributing more than 90% of the cumulative feature importance scores. Also, they (‘al’, ‘dm’, ‘hemo’, ‘sc’) have a high cosine similarity score 0.82 with the LIME feature importance explanation scores. So, these 4 features are deemed as important by the XGBoost and LIME global feature explanation importances. Hence, the number of redundant features are (24-4) = 20 for the XGBoost model. So, as per equation (1), the Interpretability score for Xgboost model would be 20/24 = 0.83.
For the Adaboost model, based on figure 13 and 16, the Cosine Similarity scores are high only when features ‘sc’ and ‘age’ are chosen. But ‘hemo’ contributes significantly for the cumulative feature scores in it’s own feature rankings. Hence, the number of important features for interpretability are 3. So, the number of redundant features would be (24-3) = 21. So equation (1) follows the Interpretability score for Adaboost = 21/24 = 0.87.
For the Random Forest model, the LIME explanation scores are similar to that of Adaboost. So, looking at figures 4 there are 4 features ‘hemo’, ‘sc’, ‘al’, ‘htn’ that contribute towards more than 90% of the cumulative scores. But in figure 16 of the LIME explanations, we see only ‘hemo’ and ‘sc’ would result in a high cosine similarity score and the aggregate cosine similarity would reduce if we include ‘al’ and ‘htn’ in the equation 3 for calculating cosine similarity. Hence, Interpretability score for Random Forest is 22/24 = 0.91.
To calculate Fidelity score in the Xgboost we are mainly calculating the external fidelity score. External Fidelity score is used to develop an explanation approach for the underlying predictive ensemble models given it’s equivalent interpretable model. The ensemble models Random Forest, Xgboost and Adaboost all rely on different Decision Trees to make their predictions. Hence, it is easier to evaluate the explainability of these ensemble models with the help of their corresponding Decision Trees. The Decision Trees could be thought as white box models capable of explaining the predictions of the black box ensemble models to the end users. The scope of using external fidelity measure is used in many existing concurrent research, specially in [68, 20]. The external fidelity would be calculated using the Precision and Recall but in context of fidelity. In this context, true features are the ones selected by the ensemble model and explanation features are the ones selected by white box Decision Tree model.
$R=\frac{\mathrm{True\ Features}\cap\mathrm{Explanation\ Features}}{\mathrm{Explanation\ Features}}$ (2)
$P=\frac{\mathrm{True\ Features}\cap\mathrm{Explanation\ Features}}{\mathrm{True\ Features}}$ (3)
TABLE X: Feature Score in Test Set as Found by the Decision Tree
Feature Name Importance Score
hemo 0.788874
al 0.114317
dm 0.049884
sc 0.034721
age 0.012204
The precision score and the recall score would be calculated using the equations (3) and (4). The F1 score as the harmonic mean of the two scores would be calculated as the final External Fidelity Score for the ensemble model.
It is seen from table 8, that only top two features are accounting for 90% of the cumulative feature scores, hence these two are our explanation features. For the true features in Xgboost in figure 14, it is evident that 4 features are contributing for the cumulative scores, hence the number of true features are 4. So, the Precision and Recall of Xgboost in terms of Fidelity are 2/4 and 2/2 respectively. Similarly we can calculate the scores for the other two ensemble models and finally calculate F1 as the measure of external fidelity. The final scores of interpretability and fidelity are populated in table 9 for all the models with their FII and FAcc values.
For the Adaboost model, true features are 3 and for the Random Forest model, it is 4. Hence, the Precision score for Adaboost is 2/3 and 2/2 respectively. For Random Forest, it would be 2/4 and 2/2 respectively.
Figure 13: Feature Importances in Test Set for Adaboost
Figure 14: Feature Importances in Test Set for Xgboost
Cosine $\text{Cosine Similarity}=\frac{\sum_{i=1}^{n}(\text{EM}_{i}\cdot\text{LIME}_{i})}{\sqrt{\sum_{i=1}^{n}(\text{EM}_{i}^{2})}\cdot\sqrt{\sum_{i=1}^{n}(\text{LIME}_{i}^{2})}}$ (4)
The other two metrics in the interpretability are the Fidelity Interpretability Index (FII) and Fidelity Accuracy Index (FAcc). These metric scores are calculated for all the 3 ensemble models: Random Forest, Adaboost and Xgboost.
@@
TABLE XI: Interpretability Measures In the Test Set of Ensemble Models
@@
AdaBoost 0.87 0.80 0.70 0.77
Random Forest 0.91 0.67 0.61 0.66
XgBoost 0.83 0.67 0.55 0.66
Figure 15: Feature Importances in Test Set for Xgboost Using LIME
Figure 16: Feature Importances in Test Set for Adaboost Using LIME
V-D Comparison with Existing Research
@@
Exisiting research on the applications of Explainable AI techniques have mostly focused on Feature Based Techniques like Shapley Additive Explanations (SHAP). To this end, study conducted by Gabriel et al is a quite comprehensive study where they employed Case Based Reasoning (CBR) approach to identify causation of CKD from case by case basis on Colombian population [21]. Similar to the work conducted [22] on investigation of XAI models on CKD from patient dataset gathered from a local hospital in Taiwan, Local Interpretable Model Agnostic Explanations (LIME) were employed to explain the predictions of the models for each cases of CKD or non-CKD. Most of the XAI feature based tools were used like Partial Dependence Plot (PDP) and Accumulated Local Effects (ALE) to validate which features contribute in which particular value range and also to analyze the inter-feature dependence in contribution scores.
@@
The novel feature of this study is the use of Example Based Techniques like CounterFactuals and Contrastive Explanation Models (CEM) to explain what changes needed to happen to each feature for a change in the outcome variable. To the best of our knowledge, no published work so far has performed similar analysis on the tabular dataset of this CKD UCI ML one.
The Fidelity Interpretability measures were also employed to compare our study with that of [20] and we found out that both the Random Forest and Adaboost model has performed considerably better than theirs in the interpretability score. Point to note here is that we have used a more thorough approach of estimating interpretability score using cosine similarity and cumulative feature scores. Also, the process of estimating Fidelity score using Precision Recall and F1 in our ensemble models is quite different than them using just the accuracy of a interpretable decision tree classifier with the accuracy of the ensemble model. The different measures to calculate interpretability and fidelity have resulted in difference of values in the FII values where our Adaboost model significantly outperforms theirs but Xgboost model has almost 23% lower score.
@@
The difference of this research with that of [20] is that this research has eradicated features like “Specific Gravity (Sg)” from our analysis beforehand after consulting with a nephrologist in Bangladesh where her feedback was, they don’t consider at any stage the Specific Gravity to be an important factor for CKD prediction. Moreover, features like Pus Cell Count (Pcc) has also been removed as this feature is mostly used for determining Acute Kidney Injury (AKI) which is not the current scope of our research. Also, we have removed features like ‘Puss Cell Volume (Pcv)’ because of it’s effectivity in identifying later stages of CKD. Features like ‘Rbc’, ‘Rbcc’, ‘Wbcc’ are removed because ‘hemo’ already provides the clinical insight on the blood cell levels required for CKD prediction. On the data pre-processing stage, the difference of our study with the concurrent ones, lie in analyzing the pattern of missingness of data, imputation through learned representations, validating the statistical significance of features with feature co-relations, employing several techniques for feature elimination, employing example based XAI algorithms, calculating interpretability with cosine similarity and calculating external fidelity to compare with concurrent research. So, to sum up, validating all the decisions through multiple measures in the pre-processing to model application to model performance analysis, direct consultation with multiple nephrologists and modifying the study with their feedback etc. are the new scope of work performed in this study different from the existing ones.
@@ VI Discussion on the Results
The primary focus of this study is to analyze the Machine Learning models decision making process in terms of feature contribution, range of feature values, peak and constant contributions, reasons of choosing a certain prediction, changes needed from feature ends for a contrasting class prediction etc. To this end, the Interpretability Fidelity measures were employed. Analyzing in terms of accuracy metrics and interpretability measures, the Adaboost model performs better than all other models specially when Fidelity Interpretability Index (FII) is considered. But the Random Forest model’s performance can’t also be ruled out as it has higher accuracy, F1 score and Interpretability value. So, from a clinical perspective both the models are presentable to the clinicians where the Random Forest model would provide a few less features as important contributors for prediction than Adaboost model.
This study focused more on the interpretability of the models with explainable AI to make the decision-making process of the model easier to understand for the general people. All the techniques like LIME, SHAP, ALE have validated the feature contributions, inter-feature dependence, feature values in CKD vs Non-CKD predictions. For this scope of work, both the feature-based and example-based explainable AI techniques were focused. To our knowledge, no research has yet implemented the example-based techniques like Counterfactuals or Contrastive Explanation Models (CEM) for the domain of using Explainable AI in Chronic Kidney Diseases. Also, the interpretability and fidelity is not measured in any concurrent research on CKD with XAI as ours. If the ML models do well in terms of interpretability, it makes the job of the researchers, clinicians and the general people easier on all ends to understand the prediction models.
@@ VII Limitations of Current Dataset and Potential Improvements
From the dataset perspective, this study is dealing with a CKD dataset containing 250 patients with CKD and 150 patients not having CKD. This dataset collected from Apollo Hospital, India is just focused on one ethnic group for this study. So naturally this dataset lacks diversity in demographic variables like ethnicity, gender and socio-economic aspects. Since the patient distribution was organized like having CKD patients occurring consecutively and similarly non-CKD patients, dataset was shuffled to make them distributed. Presence of a significant portion of missing values in a sensitive domain like CKD is also a great limitation of this dataset. As we had to do thorough research starting from the statistical persepective of first analyzing the pattern of missingness, then employ feature engineering and consultation with our first nephrologist to use the learnt representation from observed data to impute the missing values. Also, from the Pearson Correlation, it is found very strong correlation coefficients among features which was also validated through statistical measures on the Logit model. Presence of strong inter-feature correlation only makes the job of imputation harder. Moreover, there was strong positive variation in the correlation accounting for high accuracy scores. These factors have played important roles in the prediction results of the ML models. So, the limitations of the dataset and their impact on the findings are discussed in a concise list:
  • Large amount of missing values and high inter-feature correlation make it difficult to pre-process the dataset specially, creates a dilemma of keeping and discarding independent features.
  • Since the dataset lacks information in demographic variables, a thorough analysis of the demographic variations could not be performed which is important for clinical research.
  • Lacking of information in the more useful clinical investigations makes the clinical practitioners ask critical questions about specific patient findings.
VII-A Discussion with Second Nephrologist
@@
We also had our one to one session Dr. Md. Nabiul Hassan Rana in Asgar Ali Medical College, Dhaka, Bangladesh. Both of the nephrologists (Dr. Nazneen Mahmood previously) informed us that the current dataset needs more variables for a conclusive prediction of the CKD patients. Based on the class distribution in the dataset, some of the records does not really contain conclusive reasons for declaring a patient having CKD. For example, a 24 year old patient with only history of Diabetes Mellitus had been declared having CKD without any other discriminative feature values. It is unusal for a patient of 24 years to develop CKD so that’s where their suspicion arose. Moreover, another patient with high Albumin presence was declared having CKD without any other conclusive feature values. The high protein in the urea of that patient could happen for a multitude of reasons like heavy exercise, steroid, excess protein intake etc. Some of the features like ‘Pcc’ is used much later to track the progression of CKD to AKI, not to predict the cases of CKD. Features like “Sg”, “Bgr” are not important to predict the cases of CKD. The large percentage of missing data in a lot of features is also a critical aspect to this dataset. To address the dataset issues we tried to get in touch with the contributor of the dataset in Apollo Hospital, Medurai India, but to no avail. So, the research contributions made by these nephrologists are:
  • Guiding us towards the correct means of missing data imputation for the imputed data to have clinical relevance.
  • Analyzing the existing set of features and helping us in the feature elimination methods (removing ‘Sg’, ‘Sod’, ‘Pot’, ‘Pcv’, ‘Rbc’, ‘Rbcc’ from the selected features) so that selected subset of features on this dataset provide clinical significance in CKD prediction.
  • Aiding us in analyzing the potential limitations in the feature values and the diagnosis results of this CKD dataset in terms of ethnicity, proper investigations and results, associated disease, prognosis and complications etc.
  • Assisting us in collection of data from patient visits to Asgar Ali Hospital in terms of the aforementioned category of variables.
VII-B Approach to developing a new dataset